A car leaves New Orleans heading to Atlanta at a speed of 70 MPH. Thirty minutes later, another car left Atlanta heading towards New Orleans, which is about 900 miles from Atlanta, at a speed of 90 MPH. How car are they from Atlanta when they end up meeting?
Answer by checkley77(12844)

70T+90(T-.5)=900
70T+90T-45=900
160T=900+45
160T=945
T=945/160
T=5.91 HOURS AFTER THE FIRST CAR LEFT THEY WILL MEET.
70*5.91=413.7
90(5.91-.5)=90*5.41)=496.9 MILES FROM ATLANTA.
413.7+496.9=900
900~900